# On the Preservation of Input/Output Directed Graph Informativeness under Crossover

### v1

@@ 5 Actionability
As described above, informativeness characterizes how information flowing into to the system is utilized, in the sense that it characterizes how many inputs are eventually connected to an output. A symmetric concept is actionability, which characterizes the amount of informed behavior flowing out of a system, in the sense that it characterizes how many outputs are connected from an input. The levels of actionability mirror those of informativeness: A partially actionable IOD Graph has at least one path from an input to an output, a very actionable IOD Graph has a path from some input to every output, and a fully actionable IOD Graph has a path from every input to every output. On the other extreme, a non-actionable IOD Graph has no paths from inputs to outputs.
@@ 6 Discussion
One aspect of crossover in this framework is that crossover membranes are not unique; that is to say, given an input part $\Psi$ and an output part $\Omega$, there can, in general, be multiple crossover membranes which can yield a child. Moreover, these children might vary in their informativeness. This is known in the literature; Schaffer, Whitley, and Eshelman (1992) named this the “competing conventions problem,” in the sense that the internal meaning of nodes is contextual and dependent on the particular network structure, and there is nothing which forces that meaning to be consistent across networks. So two parents could have different “conventions” with regard to mappings of inputs to outputs which would undermine the effectiveness of the crossover child. Consider Figure 8, which depicts a competing conventions problem. Suppose the correct solution is a path from I1 to O1 and a path from I2 to O2. In this case, both parents solve the problem correctly while the child gets it exactly wrong. From a “competing conventions” perspective, this means that the “convention” or internal representation in nodes of these paths was not consistent/established across parents, leading to a breakdown during crossover. [^10] Note, there is also a crossover membrane which produces a child which gets the problem exactly right, namely the crossover membrane $\left\{(N1,N4),(N2,N3)\right\}$.
@@
These two observations combined suggests that the selection of the crossover membrane will be very important as we consider the implementations of in evolutionary computation. One possible algorithmic solution is to augment the crossover operation with endogenously “learned” crossover membranes or extend similarity-based methods such as Dragoni, Azzini, and Tettamanzi (2014) to IOD Graphs. We intend to pursue this in future work.
@@
Figure 10: Distribution of IOD Graph Informativeness
(a) Informativeness by Network Density
(b) An Example IOD Graph
Figure 10(a) depicts the distribution of informativeness as the degree density of the underlying IOD Graphs varies. In particular, we consider the universe of all IOD Graphs with 3 inputs, 2 outputs, 5 intermediate nodes, supposing that each input is connected to one node and each output is connected from one node, such as seen in Figure 10(b). Then Figure 10(a) depicts, for each number of intermediate node connections, what fraction are Not, Partially, Very, and Fully informative. As we can see, non-informative graphs disappear quickly as degree density increases, but, on the other hand, very informative graphs persist to a fairly high level of degree density. If edges are somehow ‘costly’ to maintain, one might expect an intermediate level of degree density, where partially and very informative graphs together dominate. [^11] Note that the number of IOD Graphs for each number of edges varies widely, in particular, exponentially. That is, while there is exactly one graph with zero edges, and exactly one graph with 25 edges, there are over 3 million IOD Graphs with 13 edges.
@@ 7 Conclusion
Here we defined Input/Output Directed Graphs to capture a broad class of networks which connect inputs to outputs. This class includes feed-forward neural networks and perceptrons, electrical circuits, municipal water systems, chemical reaction networks, and data flow networks. Various studies have applied evolutionary methods to optimizing these networks in their distinct domains. The goal of this study is to provide a strong theoretical foundation and common framework for crossover across all types of IOD Graphs and connect crossover as defined here with our proposed measure of informativeness (and actionability), which provides a measure of the connectedness of inputs to outputs in IOD Graphs. There are many directions for future work, e.g. generalized methods of constructing crossover membranes, results for IOD Graphs with nodes of different types or other parameters, generalization of informativeness to other types of graphs, and more nuanced measures of informativeness and actionability. One area of future work is a computational implementation of crossover on general IOD Graph structures. Figure 11 shows output of one such computational implementation we are developing. In this Figure, all sixteen IO partitions of a particular IOD Graph are generated. We identify contiguousness and construct crossover membranes computationally. A general IOD Graph framework such as this will support the integration of a variety of methods and applications from different domains and support general evolutionary computation on IOD Graphs.

### v3

@@ 5 Actionability
As described above, informativeness characterizes how information flowing into the system is utilized, in the sense that it characterizes how many inputs are eventually connected to an output. A symmetric concept is actionability, which characterizes the amount of informed behavior flowing out of a system, in the sense that it characterizes how many outputs are connected from an input. The levels of actionability mirror those of informativeness: A partially actionable IOD Graph has at least one path from an input to an output, a very actionable IOD Graph has a path from some input to every output, and a fully actionable IOD Graph has a path from every input to every output. On the other extreme, a non-actionable IOD Graph has no paths from inputs to outputs.
@@ 6 Discussion
One aspect of crossover in this framework is that crossover membranes are not unique; that is to say, given an input part $\Psi$ and an output part $\Omega$, there can, in general, be multiple crossover membranes which can yield a child. Moreover, these children might vary in their informativeness. This is known in the literature; Schaffer et al., (1992) named this the “competing conventions problem,” in the sense that the internal meaning of nodes is contextual and dependent on the particular network structure, and there is nothing which forces that meaning to be consistent across networks. So two parents could have different “conventions” with regard to mappings of inputs to outputs which would undermine the effectiveness of the crossover child. Consider Figure 8, which depicts a competing conventions problem. Suppose the correct solution is a path from I1 to O1 and a path from I2 to O2. In this case, both parents solve the problem correctly while the child gets it exactly wrong. From a “competing conventions” perspective, this means that the “convention” or internal representation in nodes of these paths was not consistent/established across parents, leading to a breakdown during crossover. [^10] Note, there is also a crossover membrane which produces a child which gets the problem exactly right, namely the crossover membrane $\left\{(N1,N4),(N2,N3)\right\}$.
@@
These two observations combined suggests that the selection of the crossover membrane will be very important as we consider the implementations of in evolutionary computation. One possible algorithmic solution is to augment the crossover operation with endogenously “learned” crossover membranes or extend similarity-based methods such as Dragoni et al., (2014) to IOD Graphs. We intend to pursue this in future work.
@@
Figure 10: Distribution of IOD Graph Informativeness by Degree Density
Figure 10 depicts the distribution of informativeness as the degree density of the underlying IOD Graphs varies. In particular, we consider the universe of all IOD Graphs with 3 inputs, 2 outputs, 5 intermediate nodes, supposing that each input is connected to one node and each output is connected from one node. Then Figure 10 depicts, for each number of intermediate node connections, what fraction are Not, Partially, Very, and Fully informative. As we can see, non-informative graphs disappear quickly as degree density increases, but, on the other hand, very informative graphs persist to a fairly high level of degree density. If edges are somehow ‘costly’ to maintain, one might expect an intermediate level of degree density, where partially and very informative graphs together dominate. [^11] Note that the number of IOD Graphs for each number of edges varies widely, in particular, exponentially. That is, while there is exactly one graph with zero edges, and exactly one graph with 25 edges, there are over 3 million IOD Graphs with 13 edges.
@@
6.1 Evo-Devo
Evo-devo (short for “Evolutionary Development”) is a gene-driven process in which a genome representation is translated into phenotypes, in this case, graphs, and crossover is done on the genomes, not the graphs themselves (Jacob,, 1977; Gould,, 1977). IOD Graphs are the phenotype of this process. The genomes in evo-devo, which can have any form, are typically vectors of parameters. Evo-devo is common in applications of evolutionary methods to IOD Graphs (e.g. Arifovic and Gencay,, 2001; García-Pedrajas et al.,, 2006). Some studies, such as Dragoni et al., (2014) and Uriot and Izzo, (2020), apply crossover to genomes in a network-aware method. In an approach most similar to ours, Uriot and Izzo, (2020) build a representation between hidden layers of feed-forward neural networks based on functionality and applies crossover to that representation, while our crossover allows for arbitrary crossover cuts and for a larger class of graphs.
By contrast with evo-devo, the crossover process described here is directly applied on the network topology and therefore is agnostic with respect to the genome representation. IOD Graph Crossover could lead to improved performance of evolution if the ‘direct’ representation of the graph fosters Lamarckian evolution (in which acquired traits can be inherited to offspring) through maintaining coherent sub-structures. This could be important for e.g. modifying trained neural networks.
Many existing evo-devo processes will likely not achieve offspring consistent with IOD Graph Crossover as described in Definition 7. However, if a particular evo-devo process could be shown to obey all the properties in Definition 7, then such a process would therefore follow all theorems here with regard to e.g. informativeness and actionability preservation, as well as any future theorems for IOD Graph crossover. By contrast, the crossover method in the study Uriot and Izzo, (2020), described above, does not require preserving the number of links at the crossover as we require for crossover compatibility (Definition 4, Section 3). Therefore the results in this paper do not apply.
Some studies apply crossover only to the weights of the neural networks, not the graph structure (e.g. Braun and Weisbrod,, 1993; Braun and Ragg,, 1997; Chandra et al.,, 2012). These are not examples of IOD Graph Crossover, which only considers crossover on structure, not weights.
@@ 7 Conclusion
Here we defined Input/Output Directed Graphs to capture a broad class of networks which connect inputs to outputs. This class includes feed-forward neural networks and perceptrons, electrical circuits, municipal water systems, chemical reaction networks, and data flow networks. Various studies have applied evolutionary methods to optimizing these networks in their distinct domains. The goal of this study is to provide a strong theoretical foundation and common framework for crossover across all types of IOD Graphs and connect crossover as defined here with our proposed measure of informativeness (and actionability), which provides a measure of the connectedness of inputs to outputs in IOD Graphs. There are many directions for future work, e.g. generalized methods of constructing crossover membranes, results for IOD Graphs with nodes of different types or other parameters, generalization of informativeness to other types of graphs, and more nuanced, possibly continuous, measures of informativeness and actionability. One area of future work is a computational implementation of crossover on general IOD Graph structures. Figure 11 shows output of one such computational implementation we are developing. In this Figure, all sixteen IO partitions of a particular IOD Graph are generated. We identify contiguousness and construct crossover membranes computationally. A general IOD Graph framework such as this will support the integration of a variety of methods and applications from different domains and support general evolutionary computation on IOD Graphs.